OpenStudy (anonymous):

Using complete sentences, explain how to convert your birthday function into standard form.

OpenStudy (anonymous):

function Y = (x - 11)^2 + 27

OpenStudy (anonymous):

@EMoneh97

OpenStudy (anonymous):

Standard form is A^2 + B = C right?

OpenStudy (anonymous):

standard form is y = ax^2 + bx +c

OpenStudy (anonymous):

Alight so how can I convert my function to standard?

OpenStudy (anonymous):

hey you there?

OpenStudy (anonymous):

So use FOIL on (x - 11)² and combine it with 27. Then put it in the Standard form above.

OpenStudy (anonymous):

Can you walk me through how to use FOIL on it?

OpenStudy (anonymous):

(x-11)^2 + 27 so x * 27 = 27x? then what? theres only one term on the right side

OpenStudy (anonymous):

umm could you help me step by step with mine?